Find the final F value from P = $1,500 if the i% keeps changing in the next 15 years as following:
5% for the first 4 years, 8% for the next 5 years, and 10% for the last 6 years.
Answer #1

Given P=$1500

F=P*(F/P,5%,4)*(F/P,8%,5)*(F/P,10%,6)

F=1500*(F/P,5%,4)*(F/P,8%,5)*(F/P,10%,6)

Let us find interest factors (These can be found directly from standard interest factor tables)

(F/P,5%,4)=(1+5%)^4=1.215506

(F/P,8%,5)=(1+8%)^5=1.469328

(F/P,10%,6)=(1+10%)^6=1.771561

F=1500*1.215506*1.469328*1.771561=$4745.95